An agency surveyed people worldwide, asking them, Do you live in a household with more than one car? Of the 1243 respondentPlease help how to solve this??
Answer #1

Here, , n1 = 1243 , n2 = 861
p1cap = 0.3958 , p2cap = 0.6678


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.3958 * (1-0.3958)/1243 + 0.6678*(1-0.6678)/861)
SE = 0.0212

For 0.99 CI, z-value = 2.58


MArgin of error = z *SE
= 2.58 * 0.0212
= 0.055